OpenStudy (anonymous):

Find the volume of the solid when rotated about the y-axis. y=3-x^2/9 x=3 the answer is 45/2 pi can you show me how to get that answer? thank you

OpenStudy (anonymous):

Is that x the only thing over 9 or is it the whole expression over 9?

OpenStudy (anonymous):

you are posting this again

OpenStudy (anonymous):

yea i dont get it still

OpenStudy (anonymous):

its only x^2 that is over 9

OpenStudy (anonymous):

Was it answered already @aajugdar ?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

well

OpenStudy (anonymous):

you got the points find equation with x^2 integrate it from 0 to 3

OpenStudy (anonymous):

ok x^2= 9y-27 but when i integrate it im not getting the right answer

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

Your 9y-27 is not correct.

OpenStudy (anonymous):

it should be 27-9y

OpenStudy (anonymous):

\[\int\limits_{3}^{0}\]isnt that the same thing

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

I am getting 81 pi /2

OpenStudy (anonymous):

That's the answer I got.

OpenStudy (anonymous):

am i doing this right? \[\int\limits_{3}^{0} \pi(27-9y)^2 - \pi(3)^2 \]

OpenStudy (anonymous):

(162-81)/2*pi

OpenStudy (anonymous):

I can't read that, my laptop is currently processing mass at 0% so I can't read anything from the equation editor...and is being snarky!

OpenStudy (anonymous):

math*

OpenStudy (anonymous):

hang on...trying the other method.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does that say \[y=3-\frac{1}{9}x^2?\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):